1. The sum of the ages of a mother and a daughter is 50yrs. Five years ago, the mother’s age was 7 times the age of the daughter. The present ages of mother and daughter are in year
Answer by algebrahouse.com(1659)

x = daughter's age now
50 - x = mother's age now {sum of ages is 50}

x - 5 = daughter's age five years ago {subtracted 5 from age now}
45 - x = mother's age five years ago {subtracted 5 from age now}

45 - x = 7(x - 5) {five years ago the mother was equal to 7 times the daughter}
45 - x = 7x - 35 {used distributive property}
80 = 8x {added 35 and added x to each side}
x = 10 {divided each side by 8}
50 - x = 40 {substituted 10, in for x, into 50 - x}

daughter is 10
mother is 40

For more help from me, visit: www.algebrahouse.com
Also, you can Follow @AlgebraHouse on Twitter